OpenStudy (anonymous):

solve √x+11 = (√x-5) +2

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

Solve for x: 11+sqrt(x) = sqrt(x)-3 Isolate sqrt(x) to the left hand side. Subtract 11+sqrt(x) from both sides: 0 = -14 Look for a false statement. 0 = -14 is trivially false: Answer: | | (no solutions)

OpenStudy (anonymous):

@annalizz12:First say how can we killing the powers ?!

OpenStudy (anonymous):

\[\sqrt{x+11} = \sqrt{x-5}+2\] so... \[x ^{2}+11^{2} = x ^{2} -5^{2}+2\]?

OpenStudy (anonymous):

Exacly ! Now but remeber alway the x s are in a place and numbers are in another place ! It s better !

OpenStudy (anonymous):

so where do i go from there..

OpenStudy (anonymous):

Look : we have : x+121=x+25+2 x-x=121+27 So the question has no solutions !

OpenStudy (anonymous):

ok thanks!